What is the mass percentage of hydrogen in methane?
1
Expert's answer
2020-03-05T07:57:19-0500

Answer: The mass percent of hydrogen is methane is 25 percent.

Methane Molecular weight = 16 g/mol

Formula of methane is CH4.

I mol C + 2 mol H2 = 1 mol CH4 (As, H2 is diatomic)

1 mol C =12

2 mol H2 = 4

Total mass = 4 + 12 = 16

SO, mass percentage = [Mass of hydrogen/Total mass] x 100 = 4/16 x100 = 0.25 x 100 = 25 %.